OpenStudy (anonymous):

Demond has eight bills in his wallet that are worth a total of $50. the bills are either $ 5 bills or $10 bills. How many $5 bills does desmond have?

OpenStudy (anonymous):

It could be any order for this one.@Luis Rivera if you divide by 10 that would be 1 5 dollar bill but the question isn't asking to divide by ten. :P

OpenStudy (anonymous):

So, r u saying its 1

OpenStudy (anonymous):

x+y=8 5x+10y=50

OpenStudy (anonymous):

I think that it would be 1 five dollar bill but I'm not really positive on that because it could be in any order :P

OpenStudy (anonymous):

@benice where did the 8 come from?

OpenStudy (anonymous):

you have eight bills made of x number of 5s and y number of tens

OpenStudy (anonymous):

x+y=8 5x+10y=50 hmm....how would you solve that? O.o

OpenStudy (anonymous):

x=6 y=2

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

I think benice is right. the other one would be a little confusing :P

OpenStudy (anonymous):

x+y=8 5x+10y=50 x=6 and y=2 x is 5 dollar bills y is 10 dollar bills

OpenStudy (anonymous):

nope thats not right... Check: 5(5) + 10(10) = 50 thats not right :P 25+100 doesnt equal 50 so this problem is confusing as butternots..O.o

OpenStudy (anonymous):

no wait thats right lol never mind I checked wrong!

OpenStudy (anonymous):

what is the number

OpenStudy (anonymous):

so there are 6 Five dollar bills and 2 Ten dollar bills

OpenStudy (anonymous):

10+10+5+5+5+5+5+5=50

OpenStudy (anonymous):

just one number

OpenStudy (anonymous):

Omg! Just get this: x+y=8 5x+10y=50 x=6 and y=2 x is 5 dollar bills y is 10 dollar bills ------------------------------------------------------- Just write all that down for this problem! -__-

OpenStudy (anonymous):

If you want just one number than put 6.

OpenStudy (anonymous):

there not bubbling answer i have to write in the box small, skinny boxes write it in then bubble under

OpenStudy (anonymous):

Alright so put 6. because thats how many five dollar bills there would be.

OpenStudy (anonymous):

Dude @Luis Rivera it could be MANY different combinations but due to that equation shown by @benice it makes things easier for her. More equations will confuse her more, the answer is 6. and she's done come on now :P

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

@partyrainbow276 your teacher must be on something lol she could have worded this problem in a easier way why so harsh? :P

OpenStudy (anonymous):

My teacher is a MAN i tell u a MAN not a WOMEN

OpenStudy (anonymous):

oopsie doodle :PP

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

what math is this?

OpenStudy (anonymous):

algebra or something

OpenStudy (anonymous):

My teacher is driving me CARAZY!!!!! LMFAO

OpenStudy (anonymous):

algebra or something? lolololol is it cyber school?

OpenStudy (anonymous):

Algebra is easy lol I'll tutor you on here (: